for which of the following mixtures will ag2so4(s) precipitate at 298 k

For which of the following mixtures will Ag2SO4(s) precipitate?
150.0 mL of 0.10 M Na2SO4(aq) and 5.0 mL of 0.20 M AgNO3(aq)
150.0 mL of 0.10 M Na2SO4(aq) and 5.0 mL of 0.30 M AgNO3(aq)
150.0 mL of 0.10 M Na2SO4(aq) and 5.0 mL of 0.40 M AgNO3(aq)
150.0 mL of 0.10 M Na2SO4(aq) and 5.0 mL of 0.50 M AgNO3(aq)
Answer #1
Answer #2
Concepts and reason

The solubility product for the reaction is equilibrium constant where the solid ionic compound dissociates into its ions in a solution. The solubility product is denoted asKsp{{\rm{K}}_{{\rm{sp}}}}. The solubility product value relates to the saturated solution and indicates the precipitate level of the compound. The formation precipitation starts when ionic product exceeds the solubility product.

Fundamentals

The solubility product value of the compound depends on the concentrations of its ions in a solution.

Example: AB is a solid ionic compound.

Precipitation: If the solubility product value is lesser than the concentration of the ions present in the solution, the compound precipitates in the solution.

The given solid ionic compound is Ag2SO4{\rm{A}}{{\rm{g}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_4}

The equilibrium equation for the ionic compound is given below:

Therefore, the solubility product of the ionic compound is

Ksp=[Ag+]2[SO42]=1.2×105\begin{array}{c}\\{{\rm{K}}_{{\rm{sp}}}}\,{\rm{ = }}\,{\left[ {{\rm{A}}{{\rm{g}}^{\rm{ + }}}} \right]^{\rm{2}}}\left[ {{\rm{SO}}_4^{2 - }} \right]\\\\ = 1.2 \times {10^{ - 5}}\\\end{array}

The balanced equation for the reaction of AgNO3andNa2SO4:{\rm{AgN}}{{\rm{O}}_3}\;{\rm{and N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}:

The initial concentration of [Ag+]\left[ {{\rm{A}}{{\rm{g}}^ + }} \right] ions in AgNO3{\rm{AgN}}{{\rm{O}}_3}is given below:

[Ag+]initial=5.0mL×0.20mmolAgNO3mL×1molAg+1molAgNO3(150.0+5.0)mL=0.00645M\begin{array}{c}\\{\left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}} = \frac{{5.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.20mmol AgN}}{{\rm{O}}_3}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol A}}{{\rm{g}}^ + }}}{{1{\rm{ mol AgN}}{{\rm{O}}_3}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.00645{\rm{M}}\\\end{array}

The initial concentration of [SO42]\left[ {{\rm{SO}}_4^{2 - }} \right] ions in Na2SO4{\rm{N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}is given below:

[SO42]initial=150.0mL×0.10mmolNa2SO4mL×1molSO421molNa2SO4(150.0+5.0)mL=0.0968M\begin{array}{c}\\{\left[ {{\rm{SO}}_{\rm{4}}^{{\rm{2 - }}}} \right]_{{\rm{initial}}}} = \frac{{150.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.10mmol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol }}SO_4^{2 - }}}{{1{\rm{ mol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.0968{\rm{M}}\\\end{array}

The ionic product of Ag2SO3{\rm{A}}{{\rm{g}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{3}}}is given below:

KIP=[Ag+]initial2[SO42]initial=(0.00645)2(0.0968)=4.0×106\begin{array}{c}\\{{\rm{K}}_{{\rm{IP}}}} = \left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}^2{\left[ {{\rm{SO}}_4^{2 - }} \right]_{{\rm{initial}}}}\\\\ = {\left( {0.00645} \right)^2}\left( {0.0968} \right)\\\\ = 4.0 \times {10^{ - 6}}\\\end{array}

The ionic product (KIP)\left( {{{\rm{K}}_{{\rm{IP}}}}} \right) is less than the solubility product (Ksp)\left( {{{\rm{K}}_{{\rm{sp}}}}} \right),

KIP<Kspprecipitationdoesnottakesplace{{\rm{K}}_{{\rm{IP}}}} < {{\rm{K}}_{{\rm{sp}}}} \Rightarrow {\rm{precipitation does not takes place}}

The balanced equation for the reaction of AgNO3andNa2SO4:{\rm{AgN}}{{\rm{O}}_3}\;{\rm{and N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}:

The initial concentration of [Ag+]\left[ {{\rm{A}}{{\rm{g}}^ + }} \right] ions in AgNO3{\rm{AgN}}{{\rm{O}}_3}is given below:

[Ag+]initial=5.0mL×0.30mmolAgNO3mL×1molAg+1molAgNO3(150.0+5.0)mL=0.00968M\begin{array}{c}\\{\left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}} = \frac{{5.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.30mmol AgN}}{{\rm{O}}_3}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol A}}{{\rm{g}}^ + }}}{{1{\rm{ mol AgN}}{{\rm{O}}_3}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.00968{\rm{M}}\\\end{array}

The initial concentration of [SO42]\left[ {{\rm{SO}}_4^{2 - }} \right] ions in Na2SO4{\rm{N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}is given below:

[SO42]initial=150.0mL×0.10mmolNa2SO4mL×1molSO421molNa2SO4(150.0+5.0)mL=0.0968M\begin{array}{c}\\{\left[ {{\rm{SO}}_{\rm{4}}^{{\rm{2 - }}}} \right]_{{\rm{initial}}}} = \frac{{150.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.10mmol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol }}SO_4^{2 - }}}{{1{\rm{ mol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.0968{\rm{M}}\\\end{array}

The ionic product of Ag2SO3{\rm{A}}{{\rm{g}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{3}}}is given below:

KIP=[Ag+]initial2[SO42]initial=(0.00968)2(0.0968)=9.1×106\begin{array}{c}\\{{\rm{K}}_{{\rm{IP}}}} = \left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}^2{\left[ {{\rm{SO}}_4^{2 - }} \right]_{{\rm{initial}}}}\\\\ = {\left( {0.00968} \right)^2}\left( {0.0968} \right)\\\\ = 9.1 \times {10^{ - 6}}\\\end{array}

The ionic product (KIP)\left( {{{\rm{K}}_{{\rm{IP}}}}} \right) is less than the solubility product(Ksp)\left( {{{\rm{K}}_{{\rm{sp}}}}} \right),

KIP<Kspprecipitationdoesnottakesplace{{\rm{K}}_{{\rm{IP}}}} < {{\rm{K}}_{{\rm{sp}}}} \Rightarrow {\rm{precipitation does not takes place}}

The balanced equation for the reaction of AgNO3andNa2SO4:{\rm{AgN}}{{\rm{O}}_3}\;{\rm{and N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}:

The initial concentration of [Ag+]\left[ {{\rm{A}}{{\rm{g}}^ + }} \right] ions in AgNO3{\rm{AgN}}{{\rm{O}}_3}is given below:

[Ag+]initial=5.0mL×0.40mmolAgNO3mL×1molAg+1molAgNO3(150.0+5.0)mL=0.0129M\begin{array}{c}\\{\left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}} = \frac{{5.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.40mmol AgN}}{{\rm{O}}_3}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol A}}{{\rm{g}}^ + }}}{{1{\rm{ mol AgN}}{{\rm{O}}_3}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.0129{\rm{M}}\\\end{array}

The initial concentration of [SO42]\left[ {{\rm{SO}}_4^{2 - }} \right] ions in Na2SO4{\rm{N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}is given below:

[SO42]initial=150.0mL×0.10mmolNa2SO4mL×1molSO421molNa2SO4(150.0+5.0)mL=0.0968M\begin{array}{c}\\{\left[ {{\rm{SO}}_{\rm{4}}^{{\rm{2 - }}}} \right]_{{\rm{initial}}}} = \frac{{150.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.10mmol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol }}SO_4^{2 - }}}{{1{\rm{ mol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.0968{\rm{M}}\\\end{array}

The ionic product of Ag2SO3{\rm{A}}{{\rm{g}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{3}}}is given below:

KIP=[Ag+]initial2[SO42]initial=(0.0129)2(0.0968)=1.6×105\begin{array}{c}\\{{\rm{K}}_{{\rm{IP}}}} = \left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}^2{\left[ {{\rm{SO}}_4^{2 - }} \right]_{{\rm{initial}}}}\\\\ = {\left( {0.0129} \right)^2}\left( {0.0968} \right)\\\\ = 1.6 \times {10^{ - 5}}\\\end{array}

The ionic product (KIP)\left( {{{\rm{K}}_{{\rm{IP}}}}} \right) is greater than the solubility product (Ksp)\left( {{{\rm{K}}_{{\rm{sp}}}}} \right),

KIP>Kspprecipitationtakesplace{{\rm{K}}_{{\rm{IP}}}} > {{\rm{K}}_{{\rm{sp}}}} \Rightarrow {\rm{precipitation takes place}}

The balanced equation for the reaction of AgNO3andNa2SO4:{\rm{AgN}}{{\rm{O}}_3}\;{\rm{and N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}:

The initial concentration of [Ag+]\left[ {{\rm{A}}{{\rm{g}}^ + }} \right] ions in AgNO3{\rm{AgN}}{{\rm{O}}_3}is given below:

[Ag+]initial=5.0mL×0.50mmolAgNO3mL×1molAg+1molAgNO3(150.0+5.0)mL=0.0161M\begin{array}{c}\\{\left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}} = \frac{{5.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.50mmol AgN}}{{\rm{O}}_3}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol A}}{{\rm{g}}^ + }}}{{1{\rm{ mol AgN}}{{\rm{O}}_3}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.0161{\rm{M}}\\\end{array}

The initial concentration of [SO42]\left[ {{\rm{SO}}_4^{2 - }} \right] ions in Na2SO4{\rm{N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}is given below:

[SO42]initial=150.0mL×0.10mmolNa2SO4mL×1molSO421molNa2SO4(150.0+5.0)mL=0.0968M\begin{array}{c}\\{\left[ {{\rm{SO}}_{\rm{4}}^{{\rm{2 - }}}} \right]_{{\rm{initial}}}} = \frac{{150.0{\rm{mL}} \times \frac{{{\rm{0}}{\rm{.10mmol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}{{{\rm{mL}}}} \times \frac{{1{\rm{ mol }}SO_4^{2 - }}}{{1{\rm{ mol N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}}}}}{{\left( {150.0 + 5.0} \right){\rm{mL}}}}\\\\ = 0.0968{\rm{M}}\\\end{array}

The ionic product of Ag2SO3{\rm{A}}{{\rm{g}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{3}}}is given below:

KIP=[Ag+]initial2[SO42]initial=(0.0161)2(0.0968)=2.5×105\begin{array}{c}\\{{\rm{K}}_{{\rm{IP}}}} = \left[ {{\rm{A}}{{\rm{g}}^ + }} \right]_{{\rm{initial}}}^2{\left[ {{\rm{SO}}_4^{2 - }} \right]_{{\rm{initial}}}}\\\\ = {\left( {0.0161} \right)^2}\left( {0.0968} \right)\\\\ = 2.5 \times {10^{ - 5}}\\\end{array}

The ionic product (KIP)\left( {{{\rm{K}}_{{\rm{IP}}}}} \right) is greater than the solubility product(Ksp)\left( {{{\rm{K}}_{{\rm{sp}}}}} \right),

KIP>Kspprecipitationtakesplace{{\rm{K}}_{{\rm{IP}}}} > {{\rm{K}}_{{\rm{sp}}}} \Rightarrow {\rm{precipitation takes place}}

Ans:

For the mixture, 150.0mLof0.10MNa2SO4(aq)and5.0mLof0.20MAgNO3(aq){\rm{150}}{\rm{.0mL of 0}}{\rm{.10 M N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}\left( {{\rm{aq}}} \right){\rm{ and 5}}{\rm{.0 mL of 0}}{\rm{.20 M AgN}}{{\rm{O}}_{\rm{3}}}\left( {{\rm{aq}}} \right) precipitation does not take place.

For the mixture, 150.0mLof0.10MNa2SO4(aq)and5.0mLof0.30MAgNO3(aq){\rm{150}}{\rm{.0mL of 0}}{\rm{.10 M N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}\left( {{\rm{aq}}} \right){\rm{ and 5}}{\rm{.0 mL of 0}}{\rm{.30 M AgN}}{{\rm{O}}_{\rm{3}}}\left( {{\rm{aq}}} \right) precipitation does not take place.

For the mixture, 150.0mLof0.10MNa2SO4(aq)and5.0mLof0.40MAgNO3(aq){\rm{150}}{\rm{.0mL of 0}}{\rm{.10 M N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}\left( {{\rm{aq}}} \right){\rm{ and 5}}{\rm{.0 mL of 0}}{\rm{.40 M AgN}}{{\rm{O}}_{\rm{3}}}\left( {{\rm{aq}}} \right) precipitation takes place.

For the mixture, 150.0mLof0.10MNa2SO4(aq)and5.0mLof0.50MAgNO3(aq){\rm{150}}{\rm{.0mL of 0}}{\rm{.10 M N}}{{\rm{a}}_{\rm{2}}}{\rm{S}}{{\rm{O}}_{\rm{4}}}\left( {{\rm{aq}}} \right){\rm{ and 5}}{\rm{.0 mL of 0}}{\rm{.50 M AgN}}{{\rm{O}}_{\rm{3}}}\left( {{\rm{aq}}} \right) precipitation takes place.
Answer #3

6. 9- sou /. Qxo-5 Ag, sou to ppt. Ag som tot testimoni + voy Kop= 1.8x10-5 To ppt Age soy, Ionic product (My % 60,9) should☺ As [soga] is same in all the mixture. We Can Calculate minimum Ago required to ppt Ago son for that [4, *J* 50,50] > Kopi f